Can someone out there help me with this problem? I put this in yesterday and received two different answers. One 8.5 and one 6 1/2. Which is it?
4 - (-3) + 1 /2 - 2 divided by (-2)
Thanks,
Sher

Answer by fractalier(6550)

Again, via PEMDAS we have
4 - (-3) + 1/2 - 2 / (-2) =
4 - (-3) + 1/2 + 1 =
7 + 1/2 + 1 =
8 1/2
I am not sure you have written the problem correctly...syntax is crucial in this kind of problem...